# LayeredMAPF: a decomposition of MAPF instance without compromising solvability

### v1

@@ 30       $P$ = Merge results($all\_subproblem\_paths$);
IV-C Decomposing to levels and sorting
@@
The order of solving agents is determined by whether an agent’s dependence path contains another agent’s start or target. If $a_{i}$’s dependence path contains $a_{j}$’s start, it implies that $a_{i}$ must be solved later than $a_{j}$ to ensure that $a_{j}$’s start is not occupied. Conversely, if $a_{i}$’s dependence path contains $a_{j}$’s target, $a_{i}$ must be solved before $a_{j}$ to ensure that $a_{j}$’s target is not occupied. Furthermore, the order of solving levels is determined by the order of solving agents within them.
Definition 11
@@
Definition 12
Level: a level is a strongly connected component of $\mathcal{G}_{s}$.
Intuitively, a level is a group of agents that form a loop in $\mathcal{G}_{s}$, indicating that the agents within the loop must be solved simultaneously. For example, if three agents $A$, $B$, and $C$ are in the same loop such that $A\rightarrow B\rightarrow C\rightarrow A$, it implies that $A$ must be solved before $B$, $B$ before $C$, and $C$ before $A$. Thus, $A$, $B$, and $C$ can only be solved simultaneously. Unlike clusters, where the order of solving is arbitrary, the order of solving levels is determined.
@@
Definition 14
Level ordering graph $\mathcal{G}_{l}$ is a directed graph whose nodes are levels, and edges indicate whether a level must be solved earlier than another level. If there is no edge connecting two levels, this implies that there is no explicit order to solve them, although there may be an implicit order. For example, if level $A$ has no edge connecting to level $C$, but level $A$ is connected to level $B$ and level $B$ is connected to level $C$, it implies that level $A$ must be solved earlier than level $C$.
@@
1, determine each agent’s dependence path, and available agents constrained to the agents of the current cluster;
2, obtain the solving order graph $\mathcal{G}_{s}$ from the dependence paths of the current cluster;
3, identify all strong components of $\mathcal{G}_{s}$ as levels;
4, determine the relationships between levels by examining the edges that connect them. Construct the level ordering graph $\mathcal{G}_{l}$, which represents the order of solving between levels.
5, set levels that are not later than any other levels as root levels. Then, using Breadth First Search, traverse all levels in $\mathcal{G}_{l}$, where the order of the last visit to each level represents the order of solving them;
In implementation, Tarjan’s algorithm [tarjan1972depth] is utilized to determine strongly connected components from a directed graph.
It is important to note that after completing all steps of decomposition, we ensure that the final subproblems do not compromise solvability. The entire process of decomposing into levels and sorting them is depicted in Algorithm 3.
However, we cannot guarantee that the final subproblem is the optimal decomposition. Analogous to using gradient descent to solve non-linear optimization problems, we progressively get smaller and smaller subproblems, but the final result may still not be the global optimal solution.
IV-D Solving and combining
@@
In serial MAPF methods, we only need to merge separate results into a single set. However, for the second type of MAPF method, we need to add wait actions in the resulting paths to avoid conflicts between paths from different subproblems, as shown in Algorithm 4. In the following, we define serial MAPF as MAPF$(\mathcal{C}_{\mathcal{N}},\mathcal{A},P)$, and parallel MAPF methods as MAPF$(\mathcal{C}_{\mathcal{N}},\mathcal{A})$, where $\mathcal{A}$ represents agents and $P$ represents external paths that need to be avoided.
@@
The overall process of decomposing a MAPF instance into multiple subproblems, solving the subproblems, and merging their results is illustrated in Algorithm 5.
As mentioned earlier, our decomposition method ensures the solvability of subproblems by ensuring the existence of a solution under the simplified scenario. However, our method does not guarantee that we find the optimal decomposition. While we have not provided theoretical analysis regarding this issue in this manuscript, we offer empirical analysis based on extensive testing, as detailed in Section VI.
V Results of decomposition
@@
Furthermore, we analyze how each step of the decomposition process influences the overall decomposition by evaluating changes in the decomposition rate (maximum subproblem size / total number of agents) and the number of subproblems after each step. The steps are numbered sequentially, with the first step marked as ”1” and subsequent steps indicated as ”1, 2, 3,” and so forth.
We employ a classic MAPF dataset [stern2019mapf] as MAPF instances for our analysis, comprising 24 typical maps from the dataset. These maps feature an increasing number of agents, and each number of agents is randomly selected 100 times, resulting in a total of 22,300 MAPF instances. The same instances are utilized to evaluate how decomposition influences typical MAPF methods.
@@
Figure 3: These figures show our decomposition’s performance under various of maps, interms of decomposition rate, number of subproblems time cost and memory usage. We also present scale and number of free grids (in brackets).
@@
Figure 4: These figures illustrate the performance of our decomposition method across various maps, focusing on decomposition rate, number of subproblems, time cost, and memory usage. Additionally, we provide information on the scale and number of free grids (in brackets).
V-A Decomposition rate
@@
V-A1 Performance under different maps
As depicted in Fig. 3, the final decomposition rate (i.e., the decomposition rate after step 3) demonstrates a consistent increase across almost all maps, with the exception of maps containing a high number of free grids (e.g., den520, Berlin_1_256). In these maps, the number of free grids closely aligns with the number of agents, resulting in fewer opportunities to alter dependence paths and decompose the instance. Consequently, the size of the largest subproblem approaches that of the original MAPF instance. This limitation becomes increasingly significant as the number of agents grows, ultimately leading to a decomposition rate close to 1, indicating that decomposition becomes ineffective when agents are densely packed.
Conversely, in maps with a surplus of free grids compared to the number of agents, such as den520 and Berlin_1_256, decomposition remains highly effective despite increases in the number of agents. In these scenarios, the abundance of free grids results in numerous free grid groups, facilitating the generation of dependence paths that traverse multiple agents’ start and target locations.
V-A2 How each step contribute to decomposition
@@
In maps characterized by a surplus of free grids, exceeding the number of agents (such as Berlin_1_256, Paris_1_256, and Den520d), the maximum size of clusters is reduced to merely 1 after step 1, even with hundreds of agents. This is due to the high likelihood of an agent’s initial dependence path avoiding passing through other agents’ start or target locations. Consequently, subsequent steps 2 and 3 have no further room for decomposition, resulting in similar time costs after each step in these maps.
Conversely, in maps where the number of free grids is close to the number of agents (e.g., empty_16_16, empty_32_32, random-32-32-20, and maze-32-32-4), steps 2 and 3 contribute more to decomposition than step 1 as the number of agents increases. This is because there is a lower likelihood of an agent’s initial dependence path avoiding passing through other agents’ start or target locations. Consequently, steps 2 and 3 have the opportunity to decompose subproblems by updating dependence paths.
V-B Number of subproblems
@@
Special cases arise in maps with numerous free grids, exceeding the number of agents, such as den520d and Berlin_1_256. In these instances, decomposition divides the raw instance into subproblems, each containing only one agent, resulting in the number of subproblems equaling the number of agents in the raw MAPF instance. However, it is predictable that as the number of agents increases, decomposition will become increasingly ineffective, ultimately reducing the number of subproblems to 1.
V-C Time cost and memory usage

### v2

@@ 4.2 Bipartition of clusters
1:$\mathcal{C}_{\mathcal{N}},A$, MAPF
2:$P$
3:///// / 1, decomposition of MAPF instance
4:construct $\mathcal{G}_{c}$ from $\mathcal{C}_{\mathcal{N}},A$;
5:$R$ = Decomposing to cluster($\mathcal{G}_{c}$);
6:$L=\emptyset$;
7:for cluster $r$ in $R$ do
8:     $L^{\prime}$ = Decomposing to leve and sorting($r$, $\mathcal{G}_{c}$);
9:     merge $L^{\prime}$ to $L$;
10:end for
11:$all\_subproblem\_paths$ = $\emptyset$;
12:///// / 2, solve subproblems
13:for $l$ in $L$ do
14:     if MAPF is serial then
15:         set later subproblem’s start in $\mathcal{C}_{\mathcal{N}}$ to unpassable;
16:         $current\_solution$ = MAPF($\mathcal{C}_{\mathcal{N}},l,all\_subproblem\_paths$);
17:         add $current\_solution$ to $all\_subproblem\_paths$;
18:         reset $\mathcal{C}_{\mathcal{N}}$;
19:     else
20:         set previous subproblem’s target and later subproblem’s start in $\mathcal{C}_{\mathcal{N}}$ to unpassable;
21:         $current\_solution$ = MAPF($\mathcal{C}_{\mathcal{N}},l$);
22:         add $current\_solution$ to $all\_subproblem\_paths$;
23:         reset $\mathcal{C}_{\mathcal{N}}$;
24:     end if
25:end for
26:///// / 3, merge solutions of subproblems
27:if MAPF is serial then
28:     $P$ = $all\_subproblem\_paths$;
29:else
30:     $P$ = Merge results($all\_subproblem\_paths$);
31:end if
32:return $P$;
4.3 Decomposing to levels and sorting
@@
The order of solving agents is determined by whether an agent’s dependence path contains another agent’s start or target. If $a_{i}$’s dependence path contains $a_{j}$’s start, it implies that $a_{i}$ must be solved later than $a_{j}$ to ensure that $a_{j}$’s start is not occupied (to enable $a_{i}$ and $a_{j}$ could be in different subproblem). Conversely, if $a_{i}$’s dependence path contains $a_{j}$’s target, $a_{i}$ must be solved before $a_{j}$ to ensure that $a_{j}$’s target is not occupied. Furthermore, the order of solving levels is determined by the order of solving agents within them.
Definition 13.
@@
An example of the solving order graph can be found in Fig. 12.
Figure 12: These figures illustrate the decomposition of a MAPF instance into levels in a simple scenario. Figure A shows the instance, where “$S_{i}$” and “$T_{i}$” represent the start state and target state of agent $a_{i}$.
Figure B presents the simplified connectivity graph of agents.
Figure C depicts the relevant SATs for each agent, while Figure D shows the solving order graph ($\mathcal{G}_{s}$).
Figure E displays the initial levels determined by the solving order graph, i.e., all strongly connected components in $\mathcal{G}_{s}$, along with the related level ordering graph $\mathcal{G}_{l}$.
Figure 13: These figures show the verification of four levels from the instance in Fig. 9. We check whether there is a solution for each level’s agents while assuming that the previous level’s target states and the next level’s start states are occupied. Since we ensure that each level’s agents are not related to the previous level’s target state and the next level’s start state, all levels are considered legal. The corresponding solutions for each agent are shown as dotted arrowed lines in different colors.
Essentially, a level is a strongly connected component of $\mathcal{G}_{s}$. Intuitively, a level is a group of agents that form a loop in $\mathcal{G}_{s}$, meaning that the agents within the loop must be solved simultaneously. For example, if three agents, $A$, $B$, and $C$, are in the same loop such that $A\rightarrow B\rightarrow C\rightarrow A$, it implies that $A$ must be solved before $B$, $B$ before $C$, and $C$ before $A$. Therefore, $A$, $B$, and $C$ must be solved together in parallel. Unlike clusters, where the order of solving is arbitrary, the order of solving levels is strictly determined.
@@
Definition 15.
Level ordering graph $\mathcal{G}_{l}$ is a directed graph whose nodes are levels, and edges indicate whether a level must be solved earlier than another level. If there is no edge connecting two levels, this implies that there is no explicit order to solve them, although there may be an implicit order. For example, if level $A$ has no edge connecting to level $C$, but level $A$ is connected to level $B$ (A $>$ B) and level $B$ is connected to level $C$ (B $>$ C), it implies that level $A$ must be solved earlier than level $C$.
@@
1, determine each agent’s dependence path, and available agents constrained to the agents of the current cluster (Line 1 $\sim$ 6 in Algorithm 3);
2, obtain the solving order graph $\mathcal{G}_{s}$ from the dependence paths of the current cluster(Line 7 in Algorithm 3);
3, identify all strong components of $\mathcal{G}_{s}$ as levels (Line 8 in Algorithm 3);
4, determine the relationships between levels by examining the edges that connect them. Construct the level ordering graph $\mathcal{G}_{l}$, which represents the order of solving between level (Line 9 $\sim$ 10 in Algorithm 3);
5, set levels that are not later than any other levels as root levels. Then, using Breadth First Search, traverse all levels in $\mathcal{G}_{l}$, where the order of the last visit to each level represents the order of solving them (Line 11 $\sim$ 25 in Algorithm 3).
In implementation, Tarjan’s algorithm [26] is utilized to determine strongly connected components from a directed graph.
It is important to note that after completing all steps of decomposition, we ensure that the final subproblems do not compromise legality. The entire process of decomposing into levels and sorting them is depicted in Algorithm 3. An example of decomposing a cluster into levels can be found in Fig. 12.
4.4 Solving and Combining
@@
In serial MAPF methods, we only need to merge separate results into a single set. However, for the second type of MAPF method, we need to add wait actions in the resulting paths to avoid conflicts between paths from different subproblems, as shown in Algorithm 4.
In detail, this algorithm maintains a table that records the last visited time step for each cell, in order to determine whether a subproblem’s solution collides with the solution of a previous subproblem. Starting from $t=1$, it checks each agent’s solution to see if it conflicts with the solution of a previous subproblem (Lines 1 to 20 in Algorithm 4). If a solution collides with another solution at time $t$, a wait action is added at $t$ to resolve the conflict (Lines 24 to 32 in Algorithm 4). Otherwise, the table is updated, and the algorithm checks whether there is a conflict with other solutions at time $t+1$ (Lines 33 to 39 in Algorithm 4), continuing until the last state is reached (i.e., the process is finished). Once all solutions are completed, the algorithm exits (Lines 21 to 23 in Algorithm 4).
In the following, we define serial MAPF as MAPF$(\mathcal{C}_{\mathcal{N}},\mathcal{A},P)$, and parallel MAPF methods as MAPF$(\mathcal{C}_{\mathcal{N}},\mathcal{A})$, where $\mathcal{A}$ represents agents and $P$ represents external paths that need to be avoided.
@@
The overall process of decomposing a MAPF instance into multiple subproblems (Line 1 $\sim$ 9 in Algorithm 5), solving the subproblems (Line 10 $\sim$ 23 in Algorithm 5), and merging their results (Line 24 $\sim$ 29 in Algorithm 5) is illustrated in Algorithm 5.
As mentioned earlier, our algorithm aims to minimize the loss of solvability in subproblems by introducing a legality check. Although we have not provided a theoretical analysis regarding the percentage of solvability loss, we offer empirical analysis based on extensive testing. This is detailed in the comparison between Layered LaCAM2 and raw LaCAM2 in Section 6.
4.5 Time complexity analysis
$\bullet$ Construction of connectivity graph $\mathcal{G}_{c}$
As mentioned in “Data Structures and Network Algorithms”[27], since Tarjan’s algorithm visits each node and edge only once, the time complexity of detecting connected components in a graph $G(V,E)$ is $\mathcal{O}(|V|+|E|)$.
Assuming a MAPF instance ($\mathcal{N}$ = 2), the map’s width and height are $w$ and $h$, respectively. It contains at most $w\times h$ cells and $4\times w\times h$ edges (when all cells are passable). Therefore, the time complexity to determine the $\mathcal{G}_{c}$ of an instance is $\mathcal{O}(w\times h)$. By extension, for a MAPF instance with $\mathcal{N}>2$, the time complexity to determine $\mathcal{G}_{c}$ is $\mathcal{O}(n)$, where $n$ represents the total number of states (e.g., for $\mathcal{N}$ = 2, $n$ is the area of the map; for $\mathcal{N}$ = 3, $n$ is the volume of the map).
$\bullet$ Variants of path search
As mentioned earlier, the number of nodes in $\mathcal{G}_{c}$ is $2k$, where $k$ is the number of agents. In the worst case, the number of edges in $\mathcal{G}_{c}$ is $\frac{(2k)(2k-1)}{2}$. According to Introduction to Algorithms [28], the time complexity of BFS is $\mathcal{O}(|V|+|E|)$ (where $|V|$ and $|E|$ represent the sizes of the nodes and edges, respectively). Therefore, the time complexity of $search_{p}ath_{a}gent$ and $search_{p}ath_{S}AT$ is $\mathcal{O}(k^{2})$.
$\bullet$ Decomposing to clusters
In initialization of clusters, every agents need using $search\_path\_agent$ one time. So the time complexity of $search\_path\_agent$ are $k*\mathcal{O}(k^{2})=\mathcal{O}(k^{3})$.
Considering there are $k$ nodes and at most $\frac{k(k-1)}{2}$ edges in the graph of an agent’s relevance, $\mathcal{G}_{a}$, and that the time complexity of detecting the connected components of a graph $G(V,E)$ is $\mathcal{O}(k^{2})$.
In the best case, there are $k$ clusters, and each cluster has only one agent after detecting the connected components of $\mathcal{G}_{a}$, so no further decomposition is needed. Therefore, the total time complexity of the decomposition in the best case is $\mathcal{O}(k^{3})$ (this usually happens in maps with sparse agents, such as the “warehouse-20-40-10-2-2" map and “Boston_0_256" in Section 5). However, in the worst case, further decomposition is needed, such as the bipartition of clusters and decomposition into levels (like the “random-32-32-20" map and “room-32-32-4" in Section 5).
In the bipartition of clusters, we first determine the complexity of decomposing a cluster into two clusters. Assuming there are $m$ agents in a cluster, in the worst case, it takes $m-1$ calls to $search_{p}ath_{a}gent$ to determine that one agent needs to be moved from the remaining set to the major set, and $m-1$ agents need to be moved from the remaining set to the major set. Therefore, the time complexity of splitting one cluster into two clusters is $(m-1)^{2}\times\mathcal{O}(k^{2})=\mathcal{O}(m^{2}k^{2})$.
Considering that the larger the cluster, the greater the time cost of splitting it, the worst case in the bipartition of clusters occurs when the major set always contains only one agent (i.e., maximizing the number of agents in the remaining set) until the remaining set is empty (as illustrated in Fig. 14). This requires $m-1$ splits of a cluster into two clusters, where the number of agents in the clusters is $1,2,\dots,k$. Therefore, the time complexity of the bipartition of a cluster in the worst case is: $\sum^{k}_{m=1}\mathcal{O}(m^{2}k^{2})=\mathcal{O}(\frac{k(k+1)(2k+1)}{6}(k^{2}))=\mathcal{O}(k^{5})$.
Figure 14: This figure illustrates the case that maximize the time cost of biparition of cluster.
$\bullet$ Decomposition to level and sorting
In the initialization of levels, each agent needs to use $search_{p}ath_{S}AT$ once. For a cluster with $m$ agents, the time complexity of $search_{p}ath_{S}AT$ is $m\times\mathcal{O}(k^{2})=\mathcal{O}(mk^{2})$.
Considering there are $m$ nodes and at most $m(m-1)$ edges in $\mathcal{G}_{s}$, and that in the construction of $\mathcal{G}_{l}$, we need to run Tarjan’s algorithm on $\mathcal{G}_{s}$, whose time complexity is $\mathcal{O}(|V|+|E|)$. Therefore, the time complexity for constructing $\mathcal{G}_{l}$ is $\mathcal{O}(m^{2})$.
The time complexity of level sorting is determined by a BFS search, so its complexity depends on the size of the level-ordering graph. In the worst case, each level contains only one agent, and there are $m$ levels in $\mathcal{G}_{l}$, resulting in at most $m(m-1)$ edges in $\mathcal{G}_{l}$. Since the time complexity of both DFS and BFS is $\mathcal{O}(|V|+|E|)$, the time complexity of level sorting in the worst case is $\mathcal{O}(m^{2})$.
Considering that $m\leq k$, the time complexity for decomposing a cluster with $m$ agents into levels is $\mathcal{O}(mk^{2})+\mathcal{O}(m^{2})+\mathcal{O}(m^{2})=\mathcal{O}(mk^{2})$.
So, in summary, considering the total number of agents in all clusters is $k$, the worst-case time complexity of decomposing all clusters into levels is $\mathcal{O}(k^{3})$.
$\bullet$ Combining results
As mentioned before, serial MAPF methods can take extra paths as constraints to avoid (e.g., EECBS), so no extra actions are needed to merge subproblem solutions. However, parallel MAPF methods (e.g., LaCAM) need to add wait actions to avoid conflicts between the solutions of different subproblems.
For a MAPF instance with $k$ agents, assuming there are $m$ subproblems, let the number of agents in the $i$-th subproblem be denoted as $c_{i}$, and the makespan of their solution as $s_{i}$ for $0<i<m$. In the worst case, every subproblem needs to add wait actions equal to the sum of the previous subproblem’s makespan to ensure collision-free execution (i.e., waiting until all agents from previous subproblems have arrived at their targets, except for the first subproblem, as described in the simplified scenario). Thus, the total number of added wait actions in the worst case is $\sum_{i=1}^{m-1}c_{i}\cdot s_{i-1}$. Assuming there is an upper bound on the makespan, $T$, we have: $\sum_{i=1}^{m-1}c_{i}*s_{i-1}<=\sum_{i=1}^{m-1}c_{i}*T<\sum_{i=1}^{m-1}k*T=(m-1)k*T<k^{2}T$. Therefore, the time complexity in the worst case is $\mathcal{O}(k^{2}T)$.
It is noteworthy that the makespan of solutions varies across different MAPF methods, so the time cost of merging results also differs depending on the method used.
5 Results of decomposition
@@
Furthermore, we analyze how each step of the decomposition process influences the overall decomposition by evaluating changes in the decomposition rate (maximum subproblem size / total number of agents) and the number of subproblems after each step. For simplicity, the initialization of clusters is denoted as “IC", the bipartition of clusters as “BC", and decomposing to levels and sorting as “LS".
We employ a classic MAPF dataset [29] as MAPF instances for our analysis, comprising 24 typical maps from the dataset. These maps feature an increasing number of agents, and each number of agents is randomly selected 100 times, resulting in a total of 22,300 MAPF instances. The same instances are utilized to evaluate how decomposition influences typical MAPF methods.
@@
Figure 15: These figures show our decomposition’s performance under various of maps, as number of agents increases, in terms of decomposition rate, number of subproblems time cost and memory usage. We also present scale and number of passable cells (in brackets). The map index is also used in the subsequent figures.
@@
Figure 16: These figures illustrate the performance of our decomposition method across various maps, as number of agents increases, focusing on decomposition rate, number of subproblems, time cost, and memory usage. Additionally, we provide information on the scale and number of passable cells (in brackets). The map index is also used in the subsequent figures.
5.1 Decomposition rate
@@
5.1.1 Performance under different maps
As depicted in Fig. 15, the final decomposition rate (i.e., the decomposition rate after step 3) demonstrates a consistent increase across almost all maps, with the exception of maps containing a high number of passable cells (e.g., den520, Berlin_1_256). In these maps, the number of passable cells closely aligns with the number of agents, resulting in fewer opportunities to alter dependence paths and decompose the instance. Consequently, the size of the largest subproblem approaches that of the original MAPF instance. This limitation becomes increasingly significant as the number of agents grows, ultimately leading to a decomposition rate close to 1, indicating that decomposition becomes ineffective when agents are densely packed.
Conversely, in maps with a surplus of passable cells compared to the number of agents, such as den520 and Berlin_1_256, decomposition remains highly effective despite increases in the number of agents. In these scenarios, the abundance of passable cells results in numerous free cell groups, facilitating the generation of dependence paths that no need to traverse multiple agents’ start and target locations.
5.1.2 How each step contribute to decomposition
@@
In maps characterized by a surplus of passable cells, exceeding the number of agents (such as Berlin_1_256, Paris_1_256, and Den520d), the maximum size of clusters is reduced to merely 1 after IC, even with hundreds of agents. This is due to the high likelihood of an agent’s initial dependence path avoiding passing through other agents’ start or target locations. Consequently, subsequent BC and LS have no further room for decomposition, resulting in similar time costs after each step in these maps.
Conversely, in maps where the number of passable cells is close to the number of agents (e.g., empty_16_16, empty_32_32, random-32-32-20, and maze-32-32-4), BC and LS contribute more to decomposition than IC as the number of agents increases. This is because there is a lower likelihood of an agent’s initial dependence path avoiding passing through other agents’ start or target locations. Consequently, BC and LS have the opportunity to decompose subproblems by updating dependence paths.
5.2 Number of subproblems
@@
Special cases arise in maps with numerous passable cells, exceeding the number of agents, such as den520d and Berlin_1_256. In these instances, decomposition divides the raw instance into subproblems, each containing only one agent, resulting in the number of subproblems equaling the number of agents in the raw MAPF instance. However, it is predictable that as the number of agents increases, decomposition will become increasingly ineffective, ultimately reducing the number of subproblems to 1.
5.3 Time cost and memory usage